OpenStudy (anonymous):

Please Help! MEDAL WILL BE REWARDED! Find dy/dx by implicit differentiation. x^2(y) + y^2(x) = -2

OpenStudy (amistre64):

what have you attempted?

OpenStudy (anonymous):

ive gotten to y'=(y(2x+y))/(x(x+2y))

OpenStudy (amistre64):

x^2(y) + y^2(x) = -2 x^2'(y) + x^2(y') + y^2'(x) + y^2(x') = -2' 2xy + x^2 y' + 2yx y' + y^2 = 0 x^2 y' + 2yx y' = -2xy + y^2 y' = (-2xy + y^2)/(x^2+2yx) is what i get to

OpenStudy (amistre64):

well, i dropped a negative along the way ... - y^2

OpenStudy (anonymous):

okay, so I see where I think I went wrong in this equation...yes exactly! Am I supposed to drop that there? Or should we keep it?/

OpenStudy (amistre64):

after the implicit run, the rest is just algebra manipulations. subtracting the nony' parts from left to right, and dividing off the rest; so id keep the negative parts :)

OpenStudy (amistre64):

x^2(y) + y^2(x) = -2 x^2'(y) + x^2(y') + y^2'(x) + y^2(x') = -2' 2xy + x^2 y' + 2yx y' + y^2 = 0 x^2 y' + 2yx y' = -2xy - y^2 y' = (-2xy - y^2)/(x^2+2yx)

OpenStudy (anonymous):

okay, so just for confirmation my final answer is the one above??

OpenStudy (anonymous):

Or do I need to simplify mroie?

OpenStudy (amistre64):

im not sure what the simplest simplified form would be, that is up to whoever is grading it. My teachers tend to be more concerned about showing how i get to a solution instead of some simplification or modification of it into all sorts of identical forms

OpenStudy (anonymous):

oh okay, I got you! Thank yo so much for your help, im going to review over this now and compare to my work, I tend to make the smallest mistakes!! Thank you so much!

OpenStudy (amistre64):

youre welcome, and yeah ... the calculus parts are simple enough, its the algebra that gets ya most of the time ;)